# Domain-Invariant Representation Learning of Bird Sounds

### v1

@@ I Introduction
In recent years, deep learning models have emerged as a powerful tool to process and analyze complex bioacoustic data [2]. A key source of training data for these models comes from citizen science platforms like Xeno-Canto [3], which contains over one million annotated vocalizations from more than 10,000 species, primarily birds. These citizen-led initiatives have significantly expanded the availability of labeled wildlife sound data, enabling the development of more robust and accurate deep learning models [4].
However, a challenge arises from the difference between data collected on platforms like Xeno-Canto and PAM recordings. In citizen science platforms, recordings are typically focal—where the recorder is aimed directly at the species of interest. In contrast, PAM systems passively capture sounds within natural soundscapes, leading to recordings that contain a mix of species vocalizations and background environmental noise. This difference in recording conditions creates a domain shift, complicating the ability of models trained on focal data to generalize to soundscape recordings in PAM. In practice, we require models that can perform well across diverse and potentially unseen environments.
@@
We propose a new few-shot classification benchmark based on the BirdSet [12] dataset to evaluate the generalization capabilities of models trained on Xeno-Canto’s focal recordings and tested on diverse soundscape datasets. This benchmark is designed to assess how well models can generalize across domains in challenging few-shot scenarios. We validate our proposed loss ProtoCLR on this benchmark, demonstrating its effectiveness in improving DG in bird sound classification.
@@ III-B Prototypical Contrastive Loss (ProtoCLR)
Motivated by the gradient of SupCon, we propose ProtoCLR (Prototypical Contrastive Learning of Representations), which introduces class-level centroids into contrastive learning. The centroid for each class $y$ in the batch is computed as ${c}_{y}=\frac{\displaystyle 1}{\displaystyle|C(y)|}\sum_{i\in C(y)}{z}_{i}$, where $C(y)$ is the set of indices of examples with label $y$ and $|C(y)|$ is its size. We define the ProtoCLR loss as:
@@ IV Experiments
TABLE I: Top-1 accuracy for one-shot and five-shot classification. All reported results are the average of ten runs. For our models, the results are further averaged over three checkpoints. The highest accuracies are highlighted in bold, and the second highest are underlined.
@@ IV-A Few-Shot Classification Benchmark
The BirdSet benchmark [12] comprises two tasks: multi-label classification, where each audio recording is segmented into 5-second intervals to detect the presence of one or more species (or none), and multi-class classification, where individual bird events are detected using peak detection and bambird [23], a tool for identifying bird events in audio recordings, with each event containing a single species. In line with the MetaAudio framework, we focus on the multi-class classification task to define a few-shot evaluation for assessing the generalization capabilities of pre-trained models [24]. BirdSet offers two training sets: XCL (Xeno-Canto Large) with focal recordings across nearly 10,000 species, and XCM (Xeno-Canto Medium), a specialized subset of XCL with recordings from 411 species represented in the test datasets. The validation and test datasets contain soundscape recordings. The benchmark is detailed in Table II.
Split Dataset # of recordings # of 5s examples # of classes
Train XCL 528,434 1,401,478 9,736
XCM 89,798 189,880 411
Val POW 14,911 51,394 48
Test PER 16,802 65,479 132
NES 16,117 58,251 89
UHH 3,626 12,980 27
HSN 5,460 17,940 21
NBP 24,327 76,446 51
SSW 28,403 92,514 81
SNE 19,390 64,978 56
TABLE II: BirdSet Classification Benchmark.
@@
To make predictions, we employ SimpleShot [26] by applying mean subtraction and L2-normalization to both centroids and test feature vectors. We then select the class whose centroid is closest to the test vector as the prediction. We repeat each few-shot experiment 10 times with different random seeds and report the mean and standard deviation accuracy in Table I.
@@ IV-B Reference Systems
To compare our domain-invariant pre-training approach (SupCon and ProtoCLR), we train reference systems using cross-entropy (CE) loss as a supervised baseline and SimCLR as a self-supervised contrastive baseline. Additionally, we evaluate large-scale, state-of-the-art models in bioacoustics: the encoder of BioLingual [27], an HT-SAT transformer pre-trained on AudioSet and fine-tuned using contrastive language-audio training to align animal sounds with text captions describing the class across a large collection of data including Xeno-Canto, iNaturalist, Animal Sound Archive, …etc; and the encoder of Perch [16], an EfficientNet-B1 [28] trained on Xeno-Canto for species classification, as well as taxonomic ranks genus, family, and order.
@@ IV-D Results and Discussion
Table I presents the performance of different models on one-shot and five-shot bird sound classification tasks. ProtoCLR pre-trained on XCM consistently outperforms others in both tasks, with SupCon close behind. Additionally, ProtoCLR is more computationally efficient; for one training epoch with a batch size of 256, SupCon computes 80.4B MACs, while ProtoCLR computes only 28.3B. Notably, ProtoCLR significantly outperforms CE on average (75.0 vs. 46.2). On the other hand, SimCLR performs the worst on XCM, likely due to domain shift. Incorporating unsupervised domain generalization techniques [10] could enhance SimCLR’s generalization ability.
Pre-training on the larger XCL dataset negatively impacts the performance of ProtoCLR, SupCon, and CE models. This may be due to XCL’s broader class diversity (9,736 classes) compared to XCM (411 classes) that only contains the target classes, which leads to a loss of discriminative capacity on the target classes. XCL also has a higher coefficient of variation (1.43 v 0.43) and Gini coefficient (0.62 vs 0.24), indicating greater class imbalance, which ProtoCLR may be more sensitive to than SupCon.
@@ V Conclusion
Future work could explore techniques to scale to larger class sets without compromising performance, as well as investigate domain adaptation methods, particularly source-free adaptation, which is suited for few-shot bioacoustics when source data is unavailable or too costly to retrain on.

### v3

@@ I Introduction
In recent years, deep learning models have emerged as a powerful tool to process and analyze complex bioacoustic data [2]. A key source of training data for these models comes from citizen science platforms like Xeno-Canto (XC) [3], which contains over one million annotated vocalizations from more than 10,000 species, primarily birds. These citizen-led initiatives have significantly expanded the availability of labeled wildlife sound data, enabling the development of more robust and accurate deep learning models [4].
However, a challenge arises from the difference between data collected on platforms like XC and PAM recordings. In citizen science platforms, recordings are typically focal—where the recorder is aimed directly at the species of interest. In contrast, PAM systems passively capture sounds within natural soundscapes, leading to recordings that contain a mix of species vocalizations and background environmental noise. This difference in recording conditions creates a domain shift, complicating the ability of models trained on focal data to generalize to soundscape recordings in PAM. In practice, we require models that can perform well across diverse and potentially unseen environments.
@@
We propose a new few-shot classification benchmark based on the BirdSet [12] dataset to evaluate the generalization capabilities of models trained on XC’s focal recordings and tested on diverse soundscape datasets. This benchmark is designed to assess how well models can generalize across domains in challenging few-shot scenarios. We validate our proposed loss ProtoCLR on this benchmark, demonstrating its effectiveness in improving DG in bird sound classification.
@@ III-B Prototypical Contrastive Loss (ProtoCLR)
Motivated by the gradient of SupCon and drawing inspiration from prototypical networks [11], we propose ProtoCLR (Prototypical Contrastive Learning of Representations), which introduces class-level centroids into contrastive learning. The centroid for each class $y$ in the batch is computed as ${c}_{y}=\frac{\displaystyle 1}{\displaystyle|C(y)|}\sum_{i\in C(y)}{z}_{i}$, where $C(y)$ is the set of indices of examples with label $y$ and $|C(y)|$ is its size. We define the ProtoCLR loss as:
@@ IV Experiments
Split Dataset # of recordings # of 5s examples # of classes
Train XCL 528,434 1,401,478 9,736
XCM 89,798 189,880 411
Val POW 14,911 51,394 48
Test PER 16,802 65,479 132
NES 16,117 58,251 89
UHH 3,626 12,980 27
HSN 5,460 17,940 21
NBP 24,327 76,446 51
SSW 28,403 92,514 81
SNE 19,390 64,978 56
TABLE I: BirdSet Classification Benchmark.
@@
TABLE II: Top-1 accuracy for one-shot and five-shot classification. All reported results are the average of ten runs. For our models, the results are further averaged over three checkpoints. The highest accuracies are highlighted in bold, and the second highest are underlined.
@@ IV-A Few-Shot Classification Benchmark
The BirdSet benchmark [12] comprises two tasks: multi-label classification, where each audio recording is segmented into 5-second intervals to detect the presence of one or more species (or none), and multi-class classification, where individual bird events are detected using peak detection and bambird [23], a tool for identifying bird events in audio recordings, with each event containing a single species. In line with the MetaAudio framework, we focus on the multi-class classification task to define a few-shot evaluation for assessing the generalization capabilities of pre-trained models [24]. BirdSet offers two training sets: XCL (XC Large) with focal recordings across nearly 10,000 species, and XCM (XC Medium), a specialized subset of XCL with recordings from 411 species represented in the test datasets. The validation and test datasets contain soundscape recordings. The benchmark is detailed in Table I.
@@
To make predictions, we employ SimpleShot [26] by applying mean subtraction and L2-normalization to both centroids and test feature vectors. We then select the class whose centroid is closest to the test vector as the prediction. We repeat each few-shot experiment 10 times with different random seeds and report the mean and standard deviation accuracy in Table II.
@@ IV-B Reference Systems
To compare our domain-invariant pre-training approach (SupCon and ProtoCLR), we train reference systems using cross-entropy (CE) loss as a supervised baseline and SimCLR as a self-supervised contrastive baseline. Additionally, we evaluate large-scale, state-of-the-art models in bioacoustics: BirdAVES [27], a 24-layer transformer with 1024 units, following the same architecture as HuBERT-large [28]. It is trained on bird sounds from datasets such as AudioSet, VGGSound, XC, and iNaturalist; the encoder of BioLingual [29], an HT-SAT transformer pre-trained on AudioSet and fine-tuned using contrastive language-audio training to align animal sounds with text captions describing the class across a large collection of data including XC, iNaturalist, Animal Sound Archive, …etc; and the encoder of Perch [16], an EfficientNet-B1 [30] trained on XC for species classification, as well as taxonomic ranks genus, family, and order.
@@ IV-D Results and Discussion
Table II presents the performance of different models on one-shot and five-shot bird sound classification tasks. ProtoCLR pre-trained on XCM consistently outperforms others in both tasks, with SupCon close behind. Additionally, ProtoCLR is more computationally efficient; for one training epoch with a batch size of 256, SupCon computes 80.4B MACs, while ProtoCLR computes only 28.3B. Notably, ProtoCLR significantly outperforms CE on average (75.0 vs. 46.2).
For self-supervised methods, SimCLR outperforms BirdAVES in both one-shot and five-shot learning, suggesting that invariant learning is more effective than self-prediction methods for few-shot scenarios, consistent with findings in computer vision [35]. However, among our trained models, SimCLR performs the worst on XCM, likely due to domain shift. Incorporating unsupervised domain generalization techniques [10] could enhance SimCLR’s generalization ability.
Pre-training on the larger XCL dataset negatively impacts the performance of ProtoCLR, SupCon, and CE models. This may be due to XCL’s broader class diversity (9,736 classes) compared to XCM (411 classes) that only contains the target classes, which leads to a loss of discriminative capacity on the target classes. XCL also has a higher coefficient of variation (1.43 vs 0.43) and Gini coefficient (0.62 vs 0.24), indicating greater class imbalance, which ProtoCLR may be more sensitive to than SupCon.
@@ V Conclusion
Future work could explore techniques to scale to larger class sets without compromising performance, as well as investigate domain adaptation methods, particularly source-free adaptation, which are suited for few-shot bioacoustics when source data is unavailable or too costly to retrain on.